# Position Paper: Rethinking Empirical Research in Machine Learning: Addressing Epistemic and Methodological Challenges of Experimentation

### v1

@@ Abstract
We warn against a common but incomplete understanding of empirical research in machine learning (ML) that leads to non-replicable results, makes findings unreliable, and threatens to undermine progress in the field. To overcome this alarming situation, we call for more awareness of the plurality of ways of gaining knowledge experimentally but also of some epistemic limitations. In particular, we argue most current empirical ML research is fashioned as confirmatory research while it should rather be considered exploratory.
@@ 1 The Non-Replicable ML Research Enigma
Richard Feynman (1974) described in his 1974 Caltech commencement address, Cargo Cult Science [^1] As our paper is somewhat jargon heavy, we decided to explain terminology in a glossary in the appendix, with a superscript asterisk ()(*)( ∗ ) indicating such entries., practices used by researchers to adhere to a certain way of doing things but which he believed were contrary to scientific principles. In essence, this paper warns against a similar tendency in current empirical ML and discusses potential solutions.
@@
There is ML research that does not replicate.    From an empirical scientific perspective, non-replicable research is a fundamental problem. As Karl Popper (p. 66 1959/2002) phrased it: “non-reproducible single occurrences are of no significance to science.” [^2] Reproducible here does not refer to exact computational reproducibility but generally to arriving at the same scientific conclusions, termed replicability in this paper. Consequently, ML research that does not replicate has far-reaching epistemic and practical consequences. From an epistemological point of view, it means that research results are unreliable and, to some extent, it calls into question progress in the field. In practice, it may jeopardize applied empirical researchers’ confidence in experimental results and discourage them from applying ML methods, although these novel approaches might be beneficial. For example, ML is increasingly being used in the medical domain, and this is often promising in terms of patient benefit. However, there are also examples indicating that applied researchers (are starting to) have concerns about ML being used in this high-stakes area. Consider the quite drastic warning by Dhiman et al. (2022, p. 2): “Machine learning is often portrayed as offering many advantages […]. However, these advantages have not yet materialised into patient benefit […]. Given the increasing concern about the methodological quality and risk of bias of prediction model studies [emphasis added], caution is warranted and the lack of uptake of models in medical practice is not surprising.” That is, if the ML community does not improve rigor in empirical methodological research, we think there may be a risk of a backlash against the use of ML.
@@
In general, there is a growing body of empirical evidence showing that conclusions drawn from experimental results in ML were overly optimistic at the time of publication and could not be replicated in subsequent studies. For example, Melis et al. (2018, p. 1) “arrive at the somewhat surprising conclusion that standard LSTM architectures, when properly regularized, outperform more recent models”; Henderson et al. (2018, p. 3213) found for deep reinforcement learning “that both intrinsic (e.g. random seeds, environment properties) and extrinsic sources (e.g. hyperparameters, codebases) of non-determinism can contribute to difficulties in reproducing baseline algorithms”; Christodoulou et al. (2019, p. 12) found in a systematic review “no performance benefit of machine learning over logistic regression for clinical prediction models”; Elor & Averbuch-Elor (2022, p. 1) found in their study on data balancing in classification “that balancing does not improve prediction performance for the strong” classifiers; see also Lucic et al. (2018), Riquelme et al. (2018), Raff (2019), Herrmann et al. (2020), Ferrari Dacrema et al. (2021), Marie et al. (2021), Buchka et al. (2021), Narang et al. (2021), van den Goorbergh et al. (2022), Mateus et al. (2023), McElfresh et al. (2023), or the surveys by Liao et al. (2021) and Kapoor & Narayanan (2023) for similar findings. In concrete terms, this means: there is published ML research that is “of no significance to science”, but we do not know how much!
@@
We have been warned; don’t we listen?    We are by no means the first to raise these and related issues, and the very fact that we are not the first is a matter of even graver concern. We think that empirical research in ML finds itself in a situation where practicing questionable research practices, such as state-of-the-art-hacking (SOTA-hacking; Gencoglu et al., 2019; Hullman et al., 2022), has sometimes become more rewarding than following the long line of literature warning against it. Langley wrote an editorial on Machine Learning as an Experimental Science as early as 1988, and Drummond and Hand pointed out problems with experimental method comparison in ML already in 2006. Apart from these specific examples, there is a range of literature over the last decades dealing with similar issues (e.g., Hooker, 1995; McGeoch, 2002; Johnson, 2002; Drummond, 2009; Drummond & Japkowicz, 2010; Mannarswamy & Roy, 2018; Sculley et al., 2018; Lipton & Steinhardt, 2018; Bouthillier et al., 2019; Liao et al., 2021; D’Amour et al., 2022; Raff & Farris, 2022; Lones, 2023; Trosten, 2023). Specifically relevant is the paper by Nakkiran & Belkin (2022, p. 2), in which they note a “perceived lack of legitimacy and real lack of community for good experimental science” (still) exists. If we continue not to take these warnings seriously the amount of non-replicable research will only continue to increase, as the cited very recent empirical findings indicate.
@@
In our view, the discussed literature, raising similar points, has two main shortcomings: 1) they address only specific aspects of the problem and do not provide a comprehensive picture; 2) there is a confusion of terminology. For example, Bouthillier et al. (2019) distinguish between “exploratory” and “empirical” research. Nakkiran & Belkin (2022) use the term “good experimental research” and contrast it in particular with “improving applications”. Sculley et al. (2018) talk about “empirical advancements” and “empirical analysis” that are not complemented by sufficient “empirical rigor”. And Drummond (2006) discusses ML as an “experimental science” hardly using the term “empirical” at all.
@@ 2 The Status Quo of Empirical ML
It is important to emphasize that there have been encouraging first steps in terms of empirical ML research recently. This includes the newly created publication formats Transactions on Machine Learning Research (TMLR), Journal of Data-centric Machine Learning Research (DMLR), or the NeurIPS Datasets and Benchmarks Track launched in 2021. These venues explicitly include in their scope, e.g., “reproducibility studies of previously published results or claims” (TMLR, n.d.), “systematic analyses of existing systems on novel datasets or benchmarks that yield important new insight” (DMLR, n.d.), and “systematic analyses of existing systems on novel datasets yielding important new insight” (NeurIPS, n.d.). Further examples are the I Can’t Believe It’s Not Better! (ICBINB) Workshop Series (e.g., Forde et al., 2020) and Repository of Unexpected Negative Results (ICBINB Initiative, n.d.) and efforts towards preregistration (e.g., Albanie et al., 2021) and reproducibility (e.g., Sinha et al., 2023, ). These developments, while very important, are not sufficient in our view to overcome the problems empirical ML faces. For example, while computational reproducibility may be a necessary condition, it is not a sufficient condition for replicability (e.g., Bouthillier et al., 2019). Furthermore, while the topics in the above formats cover many important aspects of empirical ML, we feel that they do not emphasize enough the importance of true replication of research, which is paramount from an empirical perspective. Most importantly, a situation in which a line of research warning us for a long time has been largely neglected will not be overcome by such practical changes alone. It also requires a change in awareness – of the importance of proper empirical ML but maybe even more of its limitations; and that there are different, equally valid types of proper empirical inquiry. We see this lack of awareness evidenced by TMLR (n.d.) itself: “TMLR emphasizes technical correctness over subjective significance, to ensure that we facilitate scientific discourse on topics that may not yet be accepted in mainstream venues [emphasis added] but may be important in the future.” This is expressed in the talk introducing TMLR [^3] TMLR - A New Open Journal For Machine Learning: https://youtu.be/Uc1r1LfJtds too. Judging by the example of other empirical sciences, this general lack of awareness of proper empirical ML is certainly the most difficult thing to overcome. Below we discuss problems we identified as symptoms of this lack.
@@
Apparently, they identify a strong bias of the ML community towards mathematical proofs (formal science perspective) and application improvements (engineering perspective), while “good experimental science” that does not focus on one of the above is not incentivized nor encouraged. Nakkiran & Belkin (2022) see this evidenced by the lack of specific subject areas, the exclusion from recent calls for papers, the lack of explicit guidelines for reviewers, and the organization of separate workshops on experimental scientific investigation at major ML conferences. In particular, reviewers “often ask for application improvements” and “for ‘theoretical justification’ for purely experimental papers” (Nakkiran & Belkin, 2022, p. 2-3). Together these factors point to a structural problem hindering the recognition and promotion of some sorts of experimental research in ML.
We completely agree with this view but think it may not immediately be clear what distinguishes improving an application from “good experimental science” at first sight.444 To avoid misunderstandings: we do consider mathematical proofs and application improvements very valuable research! As we understand it, the focus on application improvement means that much empirical/experimental research in ML focuses on developing a new method and demonstrating that it is superior to existing methods by improving some (predictive) performance metric on specific real-world benchmark datasets. “Good experimental science”, on the other hand, is not about improving performance. It is about improving understanding and knowledge of a problem, a (class of) methods, or a phenomenon. Sculley et al. (2018, p. 2) emphasize that “[e]mpirical studies [in ML] have become challenges to be ‘won’, rather than a process for developing insight and understanding. Ideally, the benefit of working with real data is to tune and examine the behavior of an algorithm under various sampling distributions, to learn about the strengths and weaknesses of the algorithms, as one would do in controlled studies.” And Rendsburg et al. (2020, p. 9) argue, “it is particularly important that our community actively attempts to understand the inherent inductive biases, strengths, and also the weaknesses of algorithms. Finding examples where an algorithm works is important – but maybe even more important is to understand under which circumstances the algorithm produces misleading results.”
@@
Problem summary.    We argue that much empirical ML research is prone to overly optimistic, unreliable, and difficult-to-refute judgments and conclusions. Many experiments in empirical ML research are based on insufficiently operationalized experimental setups, partially due to ambiguous and inconclusive conceptualizations underlying the experiments. To draw more reliable conclusions, we need more explicit, context-specific operationalizations and clearer delineations of the abstract concepts that are to be investigated. Recall that “[o]perational analysis is an excellent diagnostic tool for revealing where our knowledge is weak, in order to guide our efforts in strengthening it” (Chang, 2004, p. 147). That sometimes “good experimental research” is not encouraged enough in ML (see Problem 2) and biased experiments still occur more often than desirable (see Problem 1), exacerbates the situation considerably. The former is an excellent approach for improving insight and understanding in the sense outlined above, biased experiments tend to make this more difficult. These aspects are becoming specifically important in deep learning where the sheer complexity of today’s models, especially of foundation models, makes mathematical analysis extremely difficult. Instead, the analysis often needs to be largely experimental and thus requires thorough experimentation at the highest possible level.
@@ 3 Improving the Status Quo: More Richness in Empirical Methodological Research
For the time being, it is worth making this distinction, yet, we discuss why it is an oversimplification from an epistemic perspective in Section 4 – even more so, because we distinguish two types of exploratory empirical methodological research in the following: [^7] It is not our intention to establish a precise terminology, but we think this structure will be of assistance to the reader. insight-oriented exploratory research in contrast to method-developing exploratory research. We think insight-oriented exploratory research is what Nakkiran & Belkin (2022) mean by “good experimental research”, and what they mean by “application improvements” is a conflation of both method-developing exploratory and (supposedly) confirmatory research.
@@
More insight-oriented exploratory research.    In principle, the good thing about moving towards more insight-oriented exploratory methodological research in ML is that there are no epistemological obstacles to overcome. The neighboring field data mining and knowledge discovery clearly has an exploratory nature and is very much in the spirit of Tukey’s exploratory data analysis. There are also already some examples of influential ML research that can be considered insight-oriented and exploratory, e.g., Frankle & Carbin (2019), Belkin et al. (2019), Recht et al. (2019), Rendsburg et al. (2020), Zhang et al. (2021), or Power et al. (2021). So, rather than epistemic aspects, it is the incentives and attitudes in scientific practice towards this type of research that are an obstacle to its successful dissemination. In particular, an alleged lack of novelty and originality is often invoked, which leads to rejections. Yet, without the esteem expressed by acceptance for publication , in particular in major ML venues, there is simply little incentive to engage in exploratory ML research. More importantly, it reinforces the impression among students and young scientists that exploratory research is not an integral part of science. It is therefore necessary to stimulate, encourage, and provide opportunities to make such research visible. Nakkiran & Belkin (2022, pp. 4-5) propose to establish a special subject area within ML conferences, Experimental Science of Machine Learning, focusing on “experimental investigation into the nature of learning and learning systems.” The types of papers outlined include those with “surprising experiments, “empirical conjectures,” “refining existing phenomena,” “formalizing intuition,” and presentation of “new measurement tool[s],” all aiming to improve the understanding of ML empirically. They also provide guidelines specifically tailored to the review of this type of research.
@@
More (actual) confirmatory research.    As outlined, we believe most current empirical ML research (“application improvements”) is a mixture of method-developing exploratory research and (supposedly) confirmatory research. [^8] In a sense, this limits the potential of the former and renders the latter largely useless, with biased experiments as a result. For this reason, we add a focus on well-designed, neutral method comparison and replication studies. The scrutiny and rigor these examples of (actual) confirmatory empirical research provide are sorely needed if we are to work toward more reliable and replicable research.
@@
Neutral method comparison studies include experiments that are less biased in favor of newly proposed methods (Boulesteix et al., 2013; Lim et al., 2000; Ali & Smith, 2006; Fernández-Delgado et al., 2014). First, this includes prespecified, strictly adhered-to designs of the experimental setup, including in particular a clearly specified set of datasets and tasks. Ideally, neutral comparison studies focus on the comparison of already existing methods and are carried out by a group of authors approximately equally familiar with all the methods under consideration (Boulesteix et al., 2013). Such studies ensure more neutrality and are less prone to overly optimistic conclusions than studies proposing a method, since there is much less of an incentive to promote a particular method. Second, proper uncertainty quantification is required when analyzing empirical results in ML, especially w.r.t. the different stages of inference (e.g., model fitting, model selection, pipeline construction and performance estimation) (see Nadeau & Bengio, 2003; Bengio & Grandvalet, 2004; Hothorn et al., 2005; Bates et al., 2023). Moreover, if statistical significance testing is to be conducted to test for statistically significant performance differences across different real-world datasets, as described, e.g., by Demšar (2006), Eugster et al. (2012), Boulesteix et al. (2015a), or Eisinga et al. (2017), the methodological rigor established in other empirical domains should be applied (Munafò et al., 2017), in particular, efforts towards prior sample size calculations are important (Boulesteix et al., 2017).
@@
Moreover, we need more replication studies and meta studies. These types of research face similar reservations as insight-oriented exploratory experimental research. However, replication studies are indispensable to assess the amount of non-replicable research and to prevent it from being increased. Such studies attempt to reach the same scientific conclusions as previous studies, to provide additional empirical evidence for observed phenomena. Meta studies analyze and summarize the so accumulated evidence on a specific phenomenon. This process is the default to reach conclusions in other sciences and is important as single studies can be false and/or contradict each other. In ML, this can range from studies that attempt to replicate an experiment exactly (e.g., Lohmann et al., 2022) or slightly modify an experiment’s design (e.g., by using a different set of data in the replication of a neutral comparison study) to more comprehensive tuning and ablation studies of experiments conducted in method-developing research (e.g., Rendsburg et al., 2020; Kobak & Linderman, 2021). The latter certainly overlaps with insight-oriented exploratory research. It is important to emphasize that it is in the nature of things that a replication is not an original or novel scientific contribution in the conventional sense, and not necessarily can ‘important new insights’ be gained beyond the replication of previously observed results. Rather, it is an explicit attempt to arrive at the same results and conclusions as a previous study. The scientific relevance, which is well acknowledged in other empirical sciences such as physics or medicine, lies in gathering additional empirical evidence for a hypothesis through a successful replication. Moreover, a replication study may, but does not necessarily, also raise epistemic questions, point to experimental improvements, or provide refined concepts, especially in failed replications.
@@
More infrastructure.    To achieve this, practical limitations also need to be overcome. We require more dedicated infrastructure to make the proposed forms of research more (easily) realizable. In particular, there is a need for more and better open databases of well-curated and well-understood datasets such as OpenML (Vanschoren et al., 2013) or OpenML Benchmarking Suites (Bischl et al., 2021). Moreover, well-maintained open-source software for systematic benchmark experiments, such as the AutoML Benchmark (Gijsbers et al., 2024), HPOBench (Eggensperger et al., 2021), NAS-Bench-Suite (Mehta et al., 2022), or AlgoPerf (Dahl et al., 2023), are needed. Platforms for public leaderboards and model sharing (e.g., Hugging Face) are another important aspect, although some of these platforms are geared towards ‘horse racing’ based on predictive performance and therefore do not necessarily also provide scientific insights or interpretability. Yet, the standards and automatic nature of such platforms have the advantage that they offer concrete reference points for criticism and debate. Finally, reviewer guidelines implementing our suggestions and dedicated venues for currently hard-to-publish empirical work will allow the full potential of empirical ML to be realized (Sculley et al., 2018; Nakkiran & Belkin, 2022).

### v2

# Position: Why We Must Rethink Empirical Research in Machine Learning

@@ Abstract
We warn against a common but incomplete understanding of empirical research in machine learning that leads to non-replicable results, makes findings unreliable, and threatens to undermine progress in the field. To overcome this alarming situation, we call for more awareness of the plurality of ways of gaining knowledge experimentally but also of some epistemic limitations. In particular, we argue most current empirical machine learning research is fashioned as confirmatory research while it should rather be considered exploratory.
@@ 1 The Non-Replicable ML Research Enigma
In his Caltech commencement address “Cargo Cult Science”, [^1] As our paper contains some jargon, we have included a glossary in the appendix; asterisks ()(*)( ∗ ) in the text denote covered terms. Richard Feynman (1974) described how researchers employ practices that conflict with scientific principles to adhere to a certain way of doing things. This position paper warns against similar tendencies in empirical research in machine learning (ML) and calls for a mindset change to address methodological and epistemic challenges of experimentation.
@@
There is ML research that does not replicate.    From an empirical scientific perspective, non-replicable research is a fundamental problem. As Karl Popper (p. 66 1959/2002) phrased it: “non-reproducible single occurrences are of no significance to science.” [^2] Reproducible here does not refer to exact computational reproducibility but generally to arriving at the same scientific conclusions, termed replicability in this paper. Consequently, ML research that does not replicate has far-reaching epistemic and practical consequences. From an epistemological point of view, it means that research results are unreliable and, to some extent, it calls into question progress in the field. In practice, it may jeopardize applied empirical researchers’ confidence in experimental results and discourage them from applying ML methods, even though these novel approaches might be beneficial. For example, ML is increasingly being used in the medical domain, and this is often promising in terms of patient benefit. However, there are also examples indicating that applied researchers (are starting to) have concerns about ML being used in this high-stakes area. Consider, for example, this quite drastic warning by Dhiman et al. (2022, p. 2): “Machine learning is often portrayed as offering many advantages […]. However, these advantages have not yet materialised into patient benefit […]. Given the increasing concern about the methodological quality and risk of bias of prediction model studies [emphasis added], caution is warranted and the lack of uptake of models in medical practice is not surprising.” That is, if the ML community does not improve rigor in empirical methodological research, we think there may be a risk of a backlash against the use of ML in practice.
@@
In general, there is a growing body of empirical evidence showing that conclusions drawn from experimental results in ML were overly optimistic at the time of publication and could not be replicated in subsequent studies. For example, Melis et al. (2018, p. 1) “arrive at the somewhat surprising conclusion that standard LSTM architectures, when properly regularized, outperform more recent models”; Henderson et al. (2018, p. 3213) found for deep reinforcement learning “that both intrinsic (e.g. random seeds, environment properties) and extrinsic sources (e.g. hyperparameters, codebases) of non-determinism can contribute to difficulties in reproducing baseline algorithms”; Christodoulou et al. (2019, p. 12) found in a systematic review “no performance benefit of machine learning over logistic regression for clinical prediction models”; Elor & Averbuch-Elor (2022, p. 1) found in their study on data balancing in classification “that balancing does not improve prediction performance for the strong” classifiers; see also Lucic et al. (2018), Riquelme et al. (2018), Raff (2019), Herrmann et al. (2020), Ferrari Dacrema et al. (2021), Marie et al. (2021), Buchka et al. (2021), Narang et al. (2021), van den Goorbergh et al. (2022), Mateus et al. (2023), McElfresh et al. (2023), or the surveys by Liao et al. (2021) and Kapoor & Narayanan (2023) for similar findings. In concrete terms, there is published ML research that is, as Popper would say, of no significance to science, but we do not know how much!
@@
We have been warned; don’t we listen?    We are by no means the first to raise these and related issues, and the very fact that we are not the first is a matter of even graver concern. We think that empirical research in ML finds itself in a situation where practicing questionable research practices, such as state-of-the-art-hacking (SOTA-hacking; Gencoglu et al., 2019; Hullman et al., 2022), has sometimes become more rewarding than following the long line of literature warning against it. Langley wrote an editorial “Machine Learning as an Experimental Science” as early as 1988, and Drummond and Hand pointed out problems with experimental method comparison in ML already in 2006. Apart from these specific examples, there is a range of literature over the last decades dealing with similar issues (e.g., Hooker, 1995; McGeoch, 2002; Johnson, 2002; Drummond, 2009; Drummond & Japkowicz, 2010; Mannarswamy & Roy, 2018; Sculley et al., 2018; Lipton & Steinhardt, 2018; Bouthillier et al., 2019; Liao et al., 2021; D’Amour et al., 2022; Raff & Farris, 2022; Lones, 2023; Trosten, 2023). Specifically relevant is the paper by Nakkiran & Belkin (2022, p. 2), in which they note a “perceived lack of legitimacy and real lack of community for good experimental science” (still) exists. If we continue not to take these warnings seriously the amount of non-replicable research will only continue to increase, as the cited very recent empirical findings indicate.
@@
In our view, the discussed literature, raising similar points, has two main shortcomings: 1) they address only specific aspects of the problem and do not provide a comprehensive picture; 2) there is a confusion of terminology. For example, Bouthillier et al. (2019) distinguish between exploratory and empirical research. Nakkiran & Belkin (2022) use the term good experimental research and contrast it in particular with improving applications. Sculley et al. (2018) talk about empirical advancements and empirical analysis that are not complemented by sufficient empirical rigor. And Drummond (2006) discusses ML as an experimental science hardly using the term empirical at all.
@@ 2 The Status Quo of Empirical ML
It is important to emphasize that there have been encouraging first steps in terms of empirical ML research recently. This includes the newly created publication formats Transactions on Machine Learning Research (TMLR), Journal of Data-centric Machine Learning Research (DMLR), or the NeurIPS Datasets and Benchmarks Track launched in 2021. These venues explicitly include in their scope, e.g., “reproducibility studies of previously published results or claims” (TMLR, n.d.), “systematic analyses of existing systems on novel datasets or benchmarks that yield important new insight” (DMLR, n.d.), and “systematic analyses of existing systems on novel datasets yielding important new insight” (NeurIPS, n.d.). Further examples are the I Can’t Believe It’s Not Better! (ICBINB) workshop series (e.g., Forde et al., 2020) and Repository of Unexpected Negative Results (ICBINB Initiative, n.d.) and efforts towards preregistration (e.g., Albanie et al., 2021) and reproducibility (e.g., Sinha et al., 2023, ). These developments, while very important, are not sufficient in our view to overcome the problems empirical ML faces. For example, while computational reproducibility may be a necessary condition, it is not a sufficient condition for replicability (e.g., Bouthillier et al., 2019). Furthermore, while the topics in the above formats cover many important aspects of empirical ML, we feel that they do not emphasize enough the importance of true replication of research, which is paramount from an empirical perspective. Most importantly, a situation in which a line of research warning us for a long time has been largely neglected will not be overcome by such practical changes alone. It also requires a change in awareness – of the importance of proper empirical ML but maybe even more of its limitations; and that there are different, equally valid types of proper empirical inquiry. We see this lack of awareness evidenced by TMLR (n.d.) itself: “TMLR emphasizes technical correctness over subjective significance, to ensure that we facilitate scientific discourse on topics that may not yet be accepted in mainstream venues [emphasis added] but may be important in the future.” This is expressed in the talk introducing TMLR, too. [^3] TMLR - A New Open Journal For Machine Learning: https://youtu.be/Uc1r1LfJtds Judging by the example of other empirical sciences, this general lack of awareness of proper empirical ML is certainly the most difficult thing to overcome. Below we discuss problems we identified as symptoms of this lack.
@@
They identify a strong bias of the ML community towards mathematical proofs (formal science perspective) and application improvements (engineering perspective), while good experimental science that does not focus on one of the above is not incentivized nor encouraged. Nakkiran & Belkin (2022) see this evidenced by the lack of specific subject areas, the exclusion from recent calls for papers, the lack of explicit guidelines for reviewers, and the organization of separate workshops on experimental scientific investigation at major ML conferences. In particular, reviewers “often ask for application improvements” and “for ‘theoretical justification’ for purely experimental papers” (Nakkiran & Belkin, 2022, pp. 2–3). Together these factors point to a structural problem hindering the recognition and promotion of some sorts of experimental research in ML.
We completely agree with this view but think it may not immediately be clear what distinguishes improving an application from good experimental science at first sight.444 To avoid misunderstandings: we do consider mathematical proofs and application improvements very valuable research! As we understand it, the focus on application improvement means that much empirical/experimental research in ML focuses on developing a new method and demonstrating that it is superior to existing methods by improving some (predictive) performance metric on specific real-world benchmark datasets. Good experimental science, on the other hand, is not about improving performance. It is about improving understanding and knowledge of a problem, a (class of) methods, or a phenomenon. Sculley et al. (2018, p. 2) emphasize that “[e]mpirical studies [in ML] have become challenges to be ‘won’, rather than a process for developing insight and understanding. Ideally, the benefit of working with real data is to tune and examine the behavior of an algorithm under various sampling distributions, to learn about the strengths and weaknesses of the algorithms, as one would do in controlled studies.” And Rendsburg et al. (2020, p. 9) argue, “it is particularly important that our community actively attempts to understand the inherent inductive biases, strengths, and also the weaknesses of algorithms. Finding examples where an algorithm works is important – but maybe even more important is to understand under which circumstances the algorithm produces misleading results.”
@@
Problem summary.    We argue that much empirical ML research is prone to overly optimistic, unreliable, and difficult-to-refute judgments and conclusions. Many experiments in empirical ML research are based on insufficiently operationalized experimental setups, partially due to ambiguous and inconclusive conceptualizations underlying the experiments. To draw more reliable conclusions, we need more explicit, context-specific operationalizations and clearer delineations of the abstract concepts that are to be investigated. Recall that “[o]perational analysis is an excellent diagnostic tool for revealing where our knowledge is weak, in order to guide our efforts in strengthening it” (Chang, 2004, p. 147). That sometimes good experimental research is not encouraged enough in ML (see Problem 2) and biased experiments still occur more often than desirable (see Problem 1), exacerbates the situation considerably. The former is an excellent approach for improving insight and understanding in the sense outlined above, biased experiments tend to make this more difficult. These aspects are becoming specifically important in deep learning where the sheer complexity of today’s models, especially of foundation models, makes mathematical analysis extremely difficult. Instead, the analysis often needs to be largely experimental and thus requires thorough experimentation at the highest possible level.
@@ 3 Improving the Status Quo: More Richness in Empirical Methodological Research
For the time being, it is worth making this distinction, yet, we discuss why it is an oversimplification from an epistemic perspective in Section 4 – even more so, because we distinguish two types of exploratory empirical methodological research in the following: [^7] It is not our intention to establish a precise terminology, but we think this structure will be of assistance to the reader. insight-oriented exploratory research in contrast to method-developing exploratory research. We think insight-oriented exploratory research is what Nakkiran & Belkin (2022) mean by good experimental research, and what they mean by application improvements is a conflation of both method-developing exploratory and (supposedly) confirmatory research.
@@
More insight-oriented exploratory research.    In principle, the good thing about moving towards more insight-oriented exploratory methodological research in ML is that there are no epistemological obstacles to overcome. The neighboring field data mining and knowledge discovery clearly has an exploratory nature and is very much in the spirit of Tukey’s exploratory data analysis. There are also already some examples of influential ML research that can be considered insight-oriented and exploratory, e.g., Frankle & Carbin (2019), Belkin et al. (2019), Recht et al. (2019), Rendsburg et al. (2020), Zhang et al. (2021), or Power et al. (2021). So, rather than epistemic aspects, it is the incentives and attitudes in scientific practice towards this type of research that are an obstacle to its successful dissemination. In particular, an alleged lack of novelty and originality is often invoked, which leads to rejections. Yet, without the esteem expressed by acceptance for publication, in particular in major ML venues, there is simply little incentive to engage in exploratory ML research. More importantly, it reinforces the impression among students and young scientists that exploratory research is not an integral part of science. It is therefore necessary to stimulate, encourage, and provide opportunities to make such research visible. Nakkiran & Belkin (2022, pp. 4–5) propose to establish a special subject area within ML conferences for “Experimental Science of Machine Learning,” focusing on “experimental investigation into the nature of learning and learning systems.” The types of papers outlined include those with “surprising experiments,” “empirical conjectures,” “refining existing phenomena,” “formalizing intuition,” and presentation of “new measurement tool[s],” all aiming to improve the understanding of ML empirically. They also provide guidelines specifically tailored to the review of this type of research.
@@
More (actual) confirmatory research.    As outlined, we believe most current empirical ML research (i.e., application improvements) is a mixture of method-developing exploratory research and (supposedly) confirmatory research. [^8] In a sense, this limits the potential of the former and renders the latter largely useless, with biased experiments as a result. For this reason, we add a focus on well-designed, neutral method comparison and replication studies. The scrutiny and rigor these examples of (actual) confirmatory empirical research provide are sorely needed if we are to work toward more reliable and replicable research.
@@
Neutral method comparison studies include experiments that are less biased in favor of newly proposed methods (Boulesteix et al., 2013; Lim et al., 2000; Ali & Smith, 2006; Fernández-Delgado et al., 2014). First, this includes prespecified, strictly adhered-to designs of the experimental setup, including in particular a clearly specified set of datasets and tasks. Ideally, neutral comparison studies focus on the comparison of already existing methods and are carried out by a group of authors approximately equally familiar with all the methods under consideration (Boulesteix et al., 2013). Such studies ensure more neutrality and are less prone to overly optimistic conclusions than studies proposing a method, since there is much less of an incentive to promote a particular method. Second, proper uncertainty quantification is required when analyzing empirical results in ML, especially w.r.t. the different stages of inference (e.g., model fitting, model selection, pipeline construction, and performance estimation) (see Nadeau & Bengio, 2003; Bengio & Grandvalet, 2004; Hothorn et al., 2005; Bates et al., 2023). Moreover, if statistical significance testing is to be conducted to test for statistically significant performance differences across different real-world datasets, as described, e.g., by Demšar (2006), Eugster et al. (2012), Boulesteix et al. (2015a), or Eisinga et al. (2017), the methodological rigor established in other empirical domains should be applied (Munafò et al., 2017), in particular, efforts towards prior sample size calculations are important (Boulesteix et al., 2017).
@@
Moreover, we need more replication studies and meta-studies. These types of research face similar reservations as insight-oriented exploratory experimental research. However, replication studies are indispensable to assess the amount of non-replicable research and to prevent it from being increased. Such studies attempt to reach the same scientific conclusions as previous studies, to provide additional empirical evidence for observed phenomena. Meta-studies analyze and summarize the so accumulated evidence on a specific phenomenon. This process is the default to reach conclusions in other sciences and is important as single studies can be false and/or contradict each other. In ML, this can range from studies that attempt to replicate an experiment exactly (e.g., Lohmann et al., 2022) or slightly modify an experiment’s design (e.g., by using a different set of data in the replication of a neutral comparison study) to more comprehensive tuning and ablation studies of experiments conducted in method-developing research (e.g., Rendsburg et al., 2020; Kobak & Linderman, 2021). The latter certainly overlaps with insight-oriented exploratory research. It is important to emphasize that it is in the nature of things that a replication is not an original or novel scientific contribution in the conventional sense, and not necessarily can important new insights be gained beyond the replication of previously observed results. Rather, it is an explicit attempt to arrive at the same results and conclusions as a previous study. The scientific relevance, which is well acknowledged in other empirical sciences such as physics or medicine, lies in gathering additional empirical evidence for a hypothesis through a successful replication. Moreover, a replication study may, but does not necessarily, also raise epistemic questions, point to experimental improvements, or provide refined concepts, especially in failed replications.
@@
More infrastructure.    To achieve this, practical limitations also need to be overcome. We require more dedicated infrastructure to make the proposed forms of research more (easily) realizable. In particular, there is a need for more and better open databases of well-curated and well-understood datasets such as OpenML (Vanschoren et al., 2013) or OpenML Benchmarking Suites (Bischl et al., 2021). Moreover, well-maintained open-source software for systematic benchmark experiments, such as the AutoML Benchmark (Gijsbers et al., 2024), HPOBench (Eggensperger et al., 2021), NAS-Bench-Suite (Mehta et al., 2022), or AlgoPerf (Dahl et al., 2023), are needed. Platforms for public leaderboards and model sharing (e.g., Hugging Face) are another important aspect, although some of these platforms are geared towards horse racing based on predictive performance and therefore do not necessarily also provide scientific insights or interpretability. Yet, the standards and automatic nature of such platforms have the advantage that they offer concrete reference points for criticism and debate. Finally, reviewer guidelines implementing our suggestions and dedicated venues for currently hard-to-publish empirical work will allow the full potential of empirical ML to be realized (Sculley et al., 2018; Nakkiran & Belkin, 2022).